OpenStudy (anonymous):

Hi! Can Any1 help me with: Explain how to do/solve binary numbers. Eg. 1001, 1111, 18, 37 into 1base to base 2 and back.

OpenStudy (anonymous):

first number is in base 2 yes?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you have a 1 in the ones place. 0 in two's place. 0 in fours place and a 1 in the eights place. i am reading from right to left. so the total is 1+8=9 in base ten

OpenStudy (anonymous):

But, how would you get to such a judgement?

OpenStudy (anonymous):

second one you have 1 in the ones 1 in the twos 1 in the fours 1 in the eigths giving 1+2+4+8=15 in base ten

OpenStudy (anonymous):

how do i interpret the number 2375? this is 5 ones, 7 tens, 3 hundreds and 2 thousands in base ten yes

OpenStudy (anonymous):

Satellite73- I could use a lot of help with my practice sheet tomorrow, think u'll be able 2 help?

OpenStudy (anonymous):

in other words it is \[5\times 10^2+7\times 10^1+3\times 10^2+2\times 10^3\]

OpenStudy (anonymous):

now you are in base two so you have only zeros and ones and the places are \[2^0\] \[2^1\] \[2^2\] \[2^3\] and so on reading from right to left

OpenStudy (anonymous):

so the base two number 10010 means no ones 1 two no 4s no 8s 1 sixteen in other words 2 + 16 = 20 base ten

OpenStudy (anonymous):

now lets convert to base 2 from base 10

OpenStudy (anonymous):

Okay, so if I have the number 101, then i would say that the first 1 is in the third base, zero is in second place and the last 1 is in first base. So, then the number would be 802. Right?

OpenStudy (anonymous):

In base 2, right?

OpenStudy (anonymous):

18 we start with. the biggest power of 2 that goes in to 18 is 16 so we have a one in the 16 place. 18-16= 2 so there are no eights and no fours and one 2. and 2-2=0 so no ones. we get 10 0 1 0 meaning reading from left to right: 1 16 0 8 0 4 1 2 0 1

OpenStudy (anonymous):

is this 101 in base 2?

OpenStudy (anonymous):

Yes, I suppose

OpenStudy (anonymous):

ok so read from right to left, not left to right.

OpenStudy (anonymous):

Just an example to help clarify

OpenStudy (anonymous):

Just an example to help clarify

OpenStudy (anonymous):

you have 1 in the ones place yes?

OpenStudy (anonymous):

0 in the twos place

OpenStudy (anonymous):

and 1 in the fours place

OpenStudy (anonymous):

fours place?

OpenStudy (anonymous):

one 4, no 2s, and 1 one. gives 4+1=5 in base ten

OpenStudy (anonymous):

yes 4s place. powers of two go 1,2,4,8,16,32,64,...

OpenStudy (anonymous):

just like powers of ten go 1,10,100,1000,10000, etc

OpenStudy (anonymous):

So, you add the numbers? Oh yes, now I remeber the 2's powers.

OpenStudy (anonymous):

so the decimal number 325 means 3*100+2*10+5*1

OpenStudy (anonymous):

the base two number 1011 means 1*8+0*4+1*2+1*1

OpenStudy (anonymous):

so in base ten that number would be 8+2+1=11

OpenStudy (anonymous):

So say you wanted a number in base ten, such as 176, then would you say 1 + 10 +100, ?

OpenStudy (anonymous):

Satellite? U there?

OpenStudy (anonymous):

in base ten 176 means 1*100+ 7*10 + 1*1

OpenStudy (anonymous):

Um...left to right or right to left?

OpenStudy (anonymous):

if you want to convert this to base 2 you have to find the highest power of 2 that is in 176

OpenStudy (anonymous):

oop should be 176 = 1*100+7*10+6*1 you already know this. this is how you learn numbers to begin with. the only new thing is using a different base

OpenStudy (anonymous):

do you want to convert 176 to base 2?

OpenStudy (anonymous):

But, was my answer for 176 correct?

OpenStudy (anonymous):

you lost me. 176 is not 1 + 10 + 100. 1 + 10 + 100=111

OpenStudy (anonymous):

base ten

OpenStudy (anonymous):

So my answer was correct? And yes, please do

OpenStudy (anonymous):

So my answer was correct? And yes, please do

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if you want to convert 176 (base 10) to base 2, you have to write the powers of 2 out, and see what is the biggest one that is less than 176

OpenStudy (anonymous):

lets try an easier number. lets convert 52 to base 2

OpenStudy (anonymous):

Okay, let me taek a different number. 15. The one is in the 10 place and the five is in 1 place, so 10 +1 = 11 for base ten, right?

OpenStudy (anonymous):

i have confused you. no we write numbers in base ten as it is so 15 in base ten is just 15

OpenStudy (anonymous):

meaning 1*10 + 5*1

OpenStudy (anonymous):

you already write numbers in base ten.

OpenStudy (anonymous):

Wait- so 176 is already in base 10?

OpenStudy (anonymous):

yes of course

OpenStudy (anonymous):

all our numbers are in base ten

OpenStudy (anonymous):

lets write 52 in base 2.

OpenStudy (anonymous):

powers of 2 are 1, 2, 4, 8, 16, 32, 64, ...

OpenStudy (anonymous):

Ok- all numbers are already in base ten.....then if you get 1101, then...you convert that to base ten which is actually the normal numbers? And in terms of 52, then five is in the two place and 2 in the one place so, 2+1=3. is this correct?

OpenStudy (anonymous):

64 is too big, there is no 64 in the base ten number 52. so we start with 32 and there will be a 1 in the 32's place. (base 2) 52-32=20 now the next power of 2 below 32 is 16, and there is a 16 in twenty. so there will be a 1 in the 16s place. 20-16=4 there is no 8 in 4 so there will be a zero in the 8s place. next power is 4 and there is a 4 in 4. so there will be a 1 in the 4s place. 4-4=0, so nothing in the twos or the ones place. we get 52 base ten is 1 1 0 1 1 base 2 meaning 32 + 16 + 0 + 4 + 0 + 0 =52

OpenStudy (anonymous):

expressing numbers in base two only uses the digits 0 and1. there is no such base 2 numbers as 52. you cannot have a 5 in the two s place. you can only have a 0 or a 1 in the two s place. all base two numbers consist only of the digits 0 and 1

OpenStudy (anonymous):

Ok, I'm starting to see how to do it. Thank you!